A curve having the condition that the slope of tangent at some ponits is two times the slope of straight line joining the same point to the origin of coordinates is a

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear  pallavi

let the curve is y=f(x)

any point on the curve is (x,f(x))

 slope of tangent at this point is  = f'(x)

slope of line joinnig by origin and (x,f(x) is  = f(x) -0 /x -0

                                                              =f(x) /x

  so given  f'(x) = 2f(x) /x   .....................1

             f'(x) /f(x)  = 2/x

           ∫ f'(x) /f(x)   dx   =  ∫2/x  dx

           lnf(x)   = 2 lnx  +lnC

                      =lnCx2

            f(x)  = CX2

 Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin